OpenStudy (anonymous):

prove that 2cotx / (1 + cot^2 x) is equal to sin 2x

OpenStudy (anonymous):

please. i need answers.

OpenStudy (anonymous):

ok got it sorry it took a while

OpenStudy (anonymous):

denominator is \[1+\cot^2(x)\] if you start with \[\sin^2(x)+\cos^2(x)=1\] divide everything by \[\sin^2(x)\] you get \[1+\cot^2(x)=\frac{1}{\sin^2(x)}\] and so that is your denominator. therefore whole expression is \[2\cot(x)\times \sin^2(x)\] \[=2\frac{\cos(x)}{\sin(x)}\times \sin^2(x)\] \[=2\cos(x)\sin(x)=\sin(2x)\]

OpenStudy (anonymous):

thank u very much

OpenStudy (anonymous):

yw